Can u please help me with this question. i am so confused!!
Find the height of the right rectangular prism.
S= 108ft^2
width=4ft
length=6 ft
What i did is try to find the surface are (2b+Ph) so i did
108=2( ) <=== (i left the parenthesis blank)+20(3.2)+(4.2)
then i got, 88=2()+h
i dont think i did it right??

Answer by mananth(16946)

surface area of a rectangular prism = 2(lb+bh+lh)
l=6,b=4
108=2(6*4+4*h+6*h)
108=2(24+10h)
/2
54=24+10h
54-24=10h
30=10h
/10
h=3
height = 3ft